OpenStudy (anonymous):

6. Identify the vertex for the graph of y = 2x2 + 8x - 3.

OpenStudy (anonymous):

a. (2, 21) b. (2, 17) c. (-2, -11) d. (-2, -27)

OpenStudy (acxbox22):

the x coordinate of the vertex can be found by -b/2a from the standard form of y=ax^2+bx+c

OpenStudy (acxbox22):

then substitue the x value in to find the y value and you will have both coordinates of the vertex

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

thx @acxbox22

OpenStudy (acxbox22):

-2 is the x coordinate 2(-2)^2+8(-2)-3=8-16-3 -8-3=-11 (-2,-11) yep you are correct